OpenStudy (anonymous):

A square poster has sides measuring 2 feet less than the sides of a square sign. If the difference between their areas is 52 square feet, find the lengths of the sides of the poster and the sign.

OpenStudy (anonymous):

looks like the last one. we can do this

OpenStudy (anonymous):

\[x^2-(x-2)^2=52\] should do it, where x is the side of the larger square

OpenStudy (anonymous):

\[x^2-(x^2-4x+4)=52\] \[4x-4=52\] \[4x=56\] \[x=16\]